OpenStudy (anonymous):

f(x) = -9/x+1 and g(x) = square root x+11 find (f o g) (-7) ??

OpenStudy (anonymous):

find \[g(-7)\] then find f of the result

OpenStudy (anonymous):

\[g(-7)=\sqrt{-7+11}=\sqrt{4}=2\] now your only job is to find \[f(2)\]